The force on a wire carrying 8.75 A is a maximum of 1.28 N when placed between the pole faces of a magnet


If the pole faces are 55.5 cm in diameter, what is the approximate strength of the magnetic field?
Answer #1

F = ilB


1.28 = 8.75*0.555*B


B = 0.2635T
Answer #2

force on couductor is F = ilB

here l = 0.555m

i = 8.75 A

F = 1.28 N

so B = F/iL

B = 1.28/( 8.75 *0.555)

B = 0.263 T
Answer #3

force on a wire of lenght L in B field with currect I

F= B I L

1.28 =B * 8.75 *0.555

B=0.263